# Alon’s transmitting problem and multicolor Beck–Spencer Lemma

### v2

@@ Abstract.
The Hamming graph $H(n,q)$ is defined on the vertex set $\{1,2,\ldots,q\}^{n}$ and two vertices are adjacent if and only if they differ in precisely one coordinate. Alon [1] proved that for any sequence $v_{1},\ldots,v_{b}$ of $b=\lceil\frac{n}{2}\rceil$ vertices of $H(n,2)$, there is a vertex whose distance from $v_{i}$ is at least $b-i+1$ for all $1\leq i\leq b$. In this note, we prove that for any $q\geq 3$ and any sequence $v_{1},\ldots,v_{b}$ of $b=\lfloor(1-\frac{1}{q})n\rfloor$ vertices of $H(n,q)$, there is a vertex whose distance from $v_{i}$ is at least $b-i+1$ for all $1\leq i\leq b$.
Alon used the Beck–Spencer Lemma [3] which, in turn, was based on the floating variable method introduced by Beck and Fiala [2] who studied combinatorial discrepancies. For our proof, we extend the Beck–Spencer Lemma by using a multicolor version of the floating variable method due to Doerr and Srivastav [5].
@@ 1. Introduction
Alon posed a transmitting problem in [1] and obtained an optimal solution. This problem can be described in terms of graph burning number (see, e.g., [4]). Let $G$ be a finite graph with the vertex set $V$. For vertices $u,v\in V$ let $d(u,v)$ denote the distance between $u$ and $v$. For a non-negative integer $k$, let $\Gamma_{k}(v)$ denote the $k$-neighbors of $v$, that is, the set of vertices $u\in V$ such that $d(u,v)\leq k$. For $v_{1},v_{2},\ldots,v_{b}\in V$ we say that $(v_{1},v_{2},\ldots,v_{b})$ is a burning sequence of length $b$ if $\Gamma_{b-1}(v_{1})\cup\Gamma_{b-2}(v_{2})\cup\cdots\cup\Gamma_{0}(v_{b})=V$. The burning number of $G$, denoted by $b(G)$, is defined to be the minimum length of the burning sequences. We can think of the vertices as processors. Suppose that there is a sender outside the graph, and it sends a message to $v_{i}$ at round $i$. Each processor that receives a message at round $j$ transmits the message to all its neighbors at round $j+1$. Then $b(G)$ is the minimum number of rounds in which all the processors share the message.
@@
To prove Lemma 1 Beck and Spencer used the so-called floating variable method based on [2] where Beck and Fiala studied combinatorial discrepancies. Then Doerr and Srivastav in [5] extended the results in [2] (and many other related results concerning discrepancies) to multicolor settings. We utilize their ideas on vector-coloring to prove Lemma 3, c.f. the proof of Theorem 4.2 in [5]. So we explain a multicolor version of the floating variable method in the next section.
@@ 2. Proofs
We want to restate Lemma 3 in the form of Lemma 1. For this, we need to assign a vector to a vertex in $H(n,q)$ so that the vector behaves nicely with respect to the inner product and the Hamming distance. To this end, we use vector-coloring introduced by Doerr and Srivastav in [5]. For a $qn$-dimensional vector ${\bm{z}}=(z_{1},\ldots,z_{qn})\in{\mathbb{R}}^{qn}$, let $\bar{z}_{j}:=(z_{q(j-1)+1},z_{q(j-1)+2},\ldots,z_{qj})\in{\mathbb{R}}^{q}$, and we also write ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})$. (We write $\bar{z}$ to indicate the vector is $q$-dimensional.) For each color $i\in[q]$, we assign a vector $\bar{c}_{i}=-\frac{1}{q}\bar{1}+\bar{e}_{i}\in{\mathbb{R}}^{q}$, that is, $\bar{c}_{i}$ has $1-\frac{1}{q}$ at the $i$-th entry, and $-\frac{1}{q}$ at the other $q-1$ entries. For example, if $q=3$ then $\bar{c}_{1}=(\frac{2}{3},-\frac{1}{3},-\frac{1}{3})$, $\bar{c}_{2}=(-\frac{1}{3},\frac{2}{3},-\frac{1}{3})$, $\bar{c}_{3}=(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})$. By definition the sum of all entries of $c_{i}$ equals 00. We also have
@@
$\tilde{Q}:=\left\{\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}:0\leq\lambda_{i}\leq 1\,(i\in[q]),\,\sum_{i=1}^{q}\lambda_{i}=1\right\}.$
@@
For the second item, let $\bar{y}=\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}\in\tilde{Q}$. Noting that $\sum_{i=2}^{q}\lambda_{i}=1-\lambda_{1}$, we have
(4) $\bar{a}\cdot\bar{y}=\bar{c}_{1}\cdot\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}=(1-\tfrac{1}{q})\lambda_{1}-\tfrac{1}{q}(1-\lambda_{1})=\lambda_{1}-\tfrac{1}{q}.$
Then $-\frac{1}{q}\leq\bar{a}\cdot\bar{y}\leq 1-\frac{1}{q}$ follows from $0\leq\lambda_{1}\leq 1$. If $\bar{a}\cdot\bar{y}=1-\frac{1}{q}$, then $\lambda_{1}=1$, that is, $\bar{y}=\bar{c}_{1}$.
By the first and second items we have $|\bar{a}\cdot\bar{x}-\bar{a}\cdot\bar{y}|\leq 1$. If equality holds, then $(\bar{a}\cdot\bar{x},\bar{a}\cdot\bar{y})=(-\frac{1}{q},1-\frac{1}{q})$ or $(1-\frac{1}{q},-\frac{1}{q})$. For the former case, we have $\bar{y}=\bar{a}\in Q$ by the second item. For the latter case, we have $\lambda_{1}=0$ by (4), and the first entry of $\bar{y}$ is $\lambda_{1}(1-\frac{1}{q})+\sum_{i=2}^{q}\lambda_{i}(-\frac{1}{q})=0-\frac{1}{q}(1-\lambda_{1})=-\frac{1}{q}$. ∎
Let ${\bm{x}}=(x_{1},\ldots,x_{qn})=(\bar{x}_{1},\ldots,\bar{x}_{n})\in{\mathbb{R}}^{qn}$ be a variable vector. We describe an algorithm updating ${\bm{x}}$ step by step in such a way that at each step ${\bm{x}}\in\tilde{Q}^{n}$, and in the end ${\bm{x}}\in Q^{n}$. We call a variable $x_{k}$ floating if $x_{k}\in\{-\frac{1}{q},1-\frac{1}{q}\}$, and fixed otherwise. Once a variable is fixed, then it stays the same value and it is treated as a constant.
@@
Then ${\bm{x}}_{0}:={\mathbf{0}}$ is a solution to Eq(0), and it satisfies (C3) as well.
@@
Note that $F_{q-1}({\bm{x}}_{s})=\emptyset$ because if $q-1$ entries in $\bar{x}_{j}$ are fixed, then the remaining entry is necessarily in $\{-\frac{1}{q},1-\frac{1}{q}\}$ by (C2) and (C3).
@@
From now on, we deal with the case $|F_{q}({\bm{x}}_{s})|\leq s-1$. Since $|F_{q}({\bm{x}}_{s})|$ is non-decreasing in $s$, and $|F_{q}({\bm{x}}_{s-1})|\geq s-1$ holds when ${\bm{x}}_{s-1}$ is determined, we may assume that $|F_{q}({\bm{x}}_{s})|=s-1$. Let $f_{r}=|F_{r}({\bm{x}}_{s})|$. Then the number of fixed entries in ${\bm{x}}_{s}$ is $\sum_{r=1}^{q}rf_{r}$, and that of floating entries is $qn-\sum_{r=1}^{q}rf_{r}$.
We want to consider a system of equations consisting of only floating entries of ${\bm{x}}_{s}$. To this end, let (C1) and (C2) be equations obtained form (C1) and (C2), respectively, by replacing the coefficients corresponding to the fixed entries with zeros. Consider the following system Eq($s$) of equations:
@@
Now we have $qn-\sum_{r=1}^{q}rf_{r}$ variables, and $(n-f_{q}-1)+(n-f_{q})=2n-2f_{q}-1$ equations. Thus the dimension of solution space is $(q-2)n-\sum_{r=1}^{q}rf_{r}+2f_{q}+1\geq 2f_{q}+1\geq 1$, and we have a non-trivial solution ${\bm{y}}\neq{\mathbf{0}}$. For every $\lambda\geq 0$, $\lambda{\bm{y}}$ also satisfies Eq($q$). We extend ${\bm{y}}\in{\mathbb{R}}^{qn-\sum_{r=1}^{q}rf_{r}}$ to ${\bm{y}}\in{\mathbb{R}}^{qn}$ by inserting 00 at the positions of the fixed variables. By continuously increasing $\lambda$ from 00, we get $\lambda{\bm{y}}$ such that one of the entries becomes $-\frac{1}{q}$ or $1-\frac{1}{q}$ for the first time. Using this $\lambda>0$, let ${\bm{x}}:={\bm{x}}_{s}+\lambda{\bm{y}}$. Then ${\bm{x}}$ satisfies (C1) for $1\leq i\leq n-s$, (C2) for $1\leq j\leq n$, and (C3). Update ${\bm{x}}_{s}:={\bm{x}}$, and return to the beginning of Step(s).
@@
Similarly, for $1\leq i\leq n$, we have ${\bm{a}}_{i}\cdot{\bm{x}}_{n-i}=0$ and $F_{q}({\bm{x}}_{n-i})\geq n-i$. Thus, letting $J=[n]\setminus F_{q}({\bm{x}}_{n-i})$, we have $|J|\leq i$ and
@@
The upper bound is proved in [6]. Here we prove the lower bound. Let $m:=\lfloor(1-\tfrac{1}{q})n\rfloor$. For arbitrary $m$ vertices $v_{1},v_{2},\ldots,v_{n}\in[q]^{n}$, we show that there exists a vertex $w\in[q]^{n}$ such that $d(v_{i},w)\geq m+1-i$ for all $1\leq i\leq m$.
@@ 3.1. The case $q\geq n$
Suppose that $b(n,q-1)\geq s$ but $b(n,q)\leq s$. Then we have a burning sequence for $H(n,q)$ of length $s$, starting from, without loss of generality, $v_{1}=q{\bm{1}}$. Since $[q]^{n}\setminus\Gamma_{s-1}(v_{1})\supset[q-1]^{n}$, we need to cover $H(n,q-1)$ in less than $s$ rounds, which is impossible because $b(n,q-1)\geq s$. ∎
@@
Since $b(n,q+1)\geq b(n,q)$, it suffices to show that $b(n,n)=n+1$. By Theorem 2 we have
@@ Acknowledgments
The author thanks Noga Alon for valuable comments, in particular, for pointing out an error in the conjecture in the earlier version, and Naoki Matsumoto for stimulating discussions. The author was supported by JSPS KAKENHI Grant Number JP23K03201.

### v3

@@ Abstract.
The Hamming graph $H(n,q)$ is defined on the vertex set $\{1,2,\ldots,q\}^{n}$ and two vertices are adjacent if and only if they differ in precisely one coordinate. Alon (1992) proved that for any sequence $v_{1},\ldots,v_{b}$ of $b=\lceil\frac{n}{2}\rceil$ vertices of $H(n,2)$, there is a vertex whose distance from $v_{i}$ is at least $b-i+1$ for all $1\leq i\leq b$. In this note, we prove that for any $q\geq 3$ and any sequence $v_{1},\ldots,v_{b}$ of $b=\lfloor(1-\frac{1}{q})n\rfloor$ vertices of $H(n,q)$, there is a vertex whose distance from $v_{i}$ is at least $b-i+1$ for all $1\leq i\leq b$.
Alon used a lemma due to Beck and Spencer (1983) which, in turn, was based on the floating variable method introduced by Beck and Fiala (1981) who studied combinatorial discrepancies. For our proof, we extend the Beck–Spencer Lemma by using a multicolor version of the floating variable method due to Doerr and Srivastav (2003).
@@ 1. Introduction
Alon posed a transmitting problem in [1] and obtained an optimal solution. This problem can be described in terms of the graph burning number (see, e.g., [4]). Let $G$ be a finite graph with the vertex set $V$. For vertices $u,v\in V$ let $d(u,v)$ denote the distance between $u$ and $v$. For a non-negative integer $k$, let $\Gamma_{k}(v)$ denote the $k$-neighbors of $v$, that is, the set of vertices $u\in V$ such that $d(u,v)\leq k$. For $v_{1},v_{2},\ldots,v_{b}\in V$ we say that $(v_{1},v_{2},\ldots,v_{b})$ is a burning sequence of length $b$ if $\Gamma_{b-1}(v_{1})\cup\Gamma_{b-2}(v_{2})\cup\cdots\cup\Gamma_{0}(v_{b})=V$. The burning number of $G$, denoted by $b(G)$, is defined to be the minimum length of the burning sequences. We can think of the vertices as processors. Suppose that there is a sender outside the graph, and it sends a message to $v_{i}$ at round $i$. Each processor that receives a message at round $j$ transmits the message to all its neighbors at round $j+1$. Then $b(G)$ is the minimum number of rounds in which all the processors share the message.
@@
To prove Lemma 1, Beck and Spencer used the so-called floating variable method based on [2] where Beck and Fiala studied combinatorial discrepancies. Then Doerr and Srivastav in [5] extended the results in [2] (and many other related results concerning discrepancies) to multicolor settings. We utilize their ideas on vector-coloring to prove Lemma 3, c.f. the proof of Theorem 4.2 in [5]. So we explain a multicolor version of the floating variable method in the next section.
@@ 2. Proofs
We want to restate Lemma 3 in the form of Lemma 1. For this, we need to assign a vector to a vertex in $H(n,q)$ so that the vector behaves nicely with respect to the inner product and the Hamming distance. To this end, we use vector-coloring introduced by Doerr and Srivastav in [5]. For a $qn$-dimensional vector ${\bm{z}}=(z_{1},\ldots,z_{qn})\in{\mathbb{R}}^{qn}$, let $\bar{z}_{j}:=(z_{q(j-1)+1},z_{q(j-1)+2},\ldots,z_{qj})\in{\mathbb{R}}^{q}$, and we also write ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})$. (We write $\bar{z}$ to indicate the vector is $q$-dimensional.) For each color $i\in[q]$, we assign a vector $\bar{c}_{i}=-\frac{1}{q}\bar{1}+\bar{e}_{i}\in{\mathbb{R}}^{q}$, that is, $\bar{c}_{i}$ has $1-\frac{1}{q}$ at the $i$-th entry, and $-\frac{1}{q}$ at the other $q-1$ entries. For example, if $q=3$ then $\bar{c}_{1}=(\frac{2}{3},-\frac{1}{3},-\frac{1}{3})$, $\bar{c}_{2}=(-\frac{1}{3},\frac{2}{3},-\frac{1}{3})$, $\bar{c}_{3}=(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})$. By definition the sum of all entries of $\bar{c}_{i}$ equals 00. We also have
@@
$\tilde{Q}:=\left\{\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}:0\leq\lambda_{i}\leq 1\,(i\in[q]),\,\sum_{i=1}^{q}\lambda_{i}\leq 1\right\}.$
Note that if $\bar{x}\in\tilde{Q}$ then the number of entries of $\bar{x}$ with value $1-\frac{1}{q}$ is at most one.
@@
For the second item, let $\bar{y}=\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}\in\tilde{Q}$. Noting that $\sum_{i=2}^{q}\lambda_{i}\leq 1-\lambda_{1}$ and
$\bar{a}\cdot\bar{y}=\bar{c}_{1}\cdot\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}=(1-\tfrac{1}{q})\lambda_{1}-\tfrac{1}{q}(\lambda_{2}+\cdots+\lambda_{n}),$
we have $-\frac{1}{q}\leq\bar{a}\cdot\bar{y}\leq 1-\frac{1}{q}$, because
(4) $1-\tfrac{1}{q}\geq(1-\tfrac{1}{q})\lambda_{1}\geq\bar{a}\cdot\bar{y}\geq(1-\tfrac{1}{q})\lambda_{1}-\tfrac{1}{q}(1-\lambda_{1})=\lambda_{1}-\tfrac{1}{q}\geq-\tfrac{1}{q}.$
If $\bar{a}\cdot\bar{y}=1-\frac{1}{q}$, then $\lambda_{1}=1$, that is, $\bar{y}=\bar{c}_{1}$.
By the first and second items we have $|\bar{a}\cdot\bar{x}-\bar{a}\cdot\bar{y}|\leq 1$. If equality holds, then $(\bar{a}\cdot\bar{x},\bar{a}\cdot\bar{y})=(-\frac{1}{q},1-\frac{1}{q})$ or $(1-\frac{1}{q},-\frac{1}{q})$. For the former case, we have $\bar{y}=\bar{a}\in Q$ by the second item. For the latter case, we have $\lambda_{1}=0$ and $\sum_{i=2}^{q}\lambda_{i}=1$ by (4), and the first entry of $\bar{y}$ is $\lambda_{1}(1-\frac{1}{q})+\sum_{i=2}^{q}\lambda_{i}(-\frac{1}{q})=0-\frac{1}{q}(1-\lambda_{1})=-\frac{1}{q}$. ∎
Let ${\bm{x}}=(x_{1},\ldots,x_{qn})=(\bar{x}_{1},\ldots,\bar{x}_{n})\in{\mathbb{R}}^{qn}$ be a variable vector. We describe an algorithm updating ${\bm{x}}$ step by step in such a way that at each step ${\bm{x}}\in\tilde{Q}^{n}$, and in the end ${\bm{x}}\in Q^{n}$. We call a variable $x_{k}$ floating if $-\frac{1}{q}<x_{k}<1-\frac{1}{q}$, and fixed if $x_{k}\in\{-\frac{1}{q},1-\frac{1}{q}\}$. Once a variable is fixed, then it stays the same value and it is treated as a constant.
Before going into detail, we explain the algorithm’s geometric meaning. Let ${\mathcal{P}}$ be the polytope obtained by taking the intersection of $\tilde{Q}^{n}$ and $n$ hyperplanes $\sum_{i=1}^{q}x_{q(j-1)+i}=0$ ($1\leq j\leq n$), cf. [6]. Starting from the origin, we move along a line until we hit a facet of ${\mathcal{P}}$, and let ${\bm{x}}_{1}$ be the intersection of the line and the facet. Then, in the facet, starting from ${\bm{x}}_{1}$, we move along another line until we hit a face of ${\mathcal{P}}$, and let ${\bm{x}}_{2}$ be the intersection. We repeat this procedure so that the dimension of the faces we hit is decreasing. Finally, we stop at a face of dimension 00, and we find one of the vertices ${\bm{x}}$ of ${\mathcal{P}}$, that is, ${\bm{x}}\in Q^{n}$. The choice of lines is crucial for this algorithm to satisfy $|{\bm{a}}_{i}\cdot{\bm{x}}|<i$ for all $i$. To make the right choice, we will solve a system of equations repeatedly as described below.
@@
Then ${\bm{x}}_{0}:={\mathbf{0}}$ is a solution to Eq(0), and it satisfies (C3) as well because ${\mathbf{0}}\in{\mathbb{R}}^{q}$ is the center of $\tilde{Q}$, and ${\mathbf{0}}\in{\mathbb{R}}^{rq}$ is inside $\tilde{Q}^{n}$.
@@
We claim that $F_{q-1}({\bm{x}}_{s})=\emptyset$. Suppose, to the contrary, that $j\in F_{q-1}({\bm{x}}_{s})$. Recall that the entries of $\bar{x}_{j}$ can contain $1-\frac{1}{q}$ at most once. If $\bar{x}_{j}$ contains $1-\frac{1}{q}$, then the other $q-2$ fixed entries are $-\frac{1}{q}$. Then, by (C2), the remaining entry is $-\frac{1}{q}$, a contradiction. If $\bar{x}_{j}$ does not contain $1-\frac{1}{q}$, then all the $q-1$ fixed entries are $-\frac{1}{q}$, and the remaining entry is $1-\frac{1}{q}$, a contradiction.
@@
From now on, we deal with the case $|F_{q}({\bm{x}}_{s})|\leq s-1$. Since $|F_{q}({\bm{x}}_{s})|$ is non-decreasing in $s$, and $|F_{q}({\bm{x}}_{s-1})|\geq s-1$ holds when ${\bm{x}}_{s-1}$ is determined, we may assume that $|F_{q}({\bm{x}}_{s})|=s-1$. Let $f_{r}=|F_{r}({\bm{x}}_{s})|$. Since $f_{q-1}=0$ and $\sum_{r=0}^{q-2}f_{r}=n-f_{q}$, it follows that $\sum_{r=1}^{q-2}rf_{r}\leq(q-2)\sum_{r=1}^{q-2}f_{r}=(q-2)(n-f_{q})$.
Recall that we treat a fixed variable as a constant. Let $[qn]=I_{c}\cup I_{v}$ be a partition, where $I_{c}$ is the set of indices for constant entries of ${\bm{x}}_{s}=(x_{1},x_{2},\ldots,x_{qn})$, and $I_{v}$ is the set of indices for floating variables of ${\bm{x}}_{s}$. Then $|I_{v}|=qn-\sum_{r=1}^{q}rf_{r}$.
Consider the following system Eq($s$) of equations:
@@
We have $|I_{v}|$ variables, and $(n-f_{q}-1)+(n-f_{q})=2n-2f_{q}-1$ equations. Thus the dimension of the solution space is
$|I_{v}|-(2n-2f_{q}-1)=(q-2)(n-f_{q})-\sum_{r=1}^{q-2}rf_{r}+1\geq 1,$
and we obtain a line of solution ${\bm{x}}_{s}+\lambda{\bm{y}}$, where $\lambda\in{\mathbb{R}}$ and $y_{i}=0$ for $i\in I_{c}$. By continuously increasing $\lambda$ from 00, we get ${\bm{x}}_{s}+\lambda{\bm{y}}$ such that one of the entries in $I_{v}$ becomes $-\frac{1}{q}$ or $1-\frac{1}{q}$ for the first time. Using this $\lambda>0$, let ${\bm{x}}:={\bm{x}}_{s}+\lambda{\bm{y}}$. Then ${\bm{x}}$ satisfies (C1) for $1\leq i\leq n-s$, (C2) for $1\leq j\leq n$, and (C3). Update ${\bm{x}}_{s}:={\bm{x}}$, and return to the beginning of Step(s).
@@
Similarly, for $1\leq i\leq n$, we have ${\bm{a}}_{i}\cdot{\bm{x}}_{n-i}=0$ and $|F_{q}({\bm{x}}_{n-i})|\geq n-i$. Thus, letting $J=[n]\setminus F_{q}({\bm{x}}_{n-i})$, we have $|J|\leq i$ and
@@
As stated in the discussion following the statement of Theorem 2, the upper bound is proved in [7]. Here we prove the lower bound. Let $m:=\lfloor(1-\tfrac{1}{q})n\rfloor$. For arbitrary $m$ vertices $v_{1},v_{2},\ldots,v_{m}\in[q]^{n}$, we show that there exists a vertex $w\in[q]^{n}$ such that $d(v_{i},w)\geq m+1-i$ for all $1\leq i\leq m$.
@@ 3.1. The case $q\geq n$
Suppose that $b(n,q-1)\geq s$ but $b(n,q)\leq s$. Then we have a burning sequence for $H(n,q)$ of length $s$, starting from, without loss of generality, $v_{1}=q\mathbf{1}$. Since $[q]^{n}\setminus\Gamma_{s-1}(v_{1})\supset[q-1]^{n}$, we need to cover $H(n,q-1)$ in less than $s$ rounds, which is impossible because $b(n,q-1)\geq s$. ∎
@@
Since $b(n,q+1)\geq b(n,q)$, it suffices to show that $b(n,n)=n+1$. This is true if $n=3$. Indeed we have $b(3,2)=3$ by Theorem 1, and so $b(3,3)\geq 4$ by Claim 3.
Now let $n\geq 4$. By Theorem 2 we have
@@ Acknowledgments
The author thanks Noga Alon for valuable comments, in particular, for pointing out an error in the conjecture in the earlier version, Hajime Tanaka for providing information on eigenpolytopes and [6], and Naoki Matsumoto for stimulating discussions. He also thanks the referees who read the manuscript very carefully and corrected some errors.
The author was supported by JSPS KAKENHI Grant Number JP23K03201.